OpenStudy (anonymous):

3. textbook company keeps track of its daily cost for printing whole textbooks to different schools but can only print up to 500 books a day. The daily cost (C) to print x number of textbooks is given by C(x) = 3.50x + 1200. 1. Using the function, determine one possible daily cost by evaluating the function for any input of your choice. 2. For what value of x does C(x) = 1900? 3. Interpret the statement C(20) = 1270. 4. Interpret the statement C(w) = p. 5. Explain whether C(40.5) = 1341.75 is correct and possible given the context of this problem.

OpenStudy (anonymous):

1.) 2950. I set X to be 500 and C(500)=2950. 2.)1900=3.50X+1200. 700=3.50X. X=200 3.) Set X as 20. [3.5(20)+1200=1270] 4.) By replacing X with w, the value of C(w)=3.50w+1200=p 5.) The question states 'whole books'. The equation may not be valid for non integer values so I'd probably say no to that one. Hope this helps